# DN-Splatter: Depth and Normal Priors for Gaussian Splatting and Meshing

### v2

@@ Abstract
High-fidelity 3D reconstruction of common indoor scenes is crucial for VR and AR applications. 3D Gaussian splatting, a novel differentiable rendering technique, has achieved state-of-the-art novel view synthesis results with high rendering speeds and relatively low training times. However, its performance on scenes commonly seen in indoor datasets is poor due to the lack of geometric constraints during optimization. We extend 3D Gaussian splatting with depth and normal cues to tackle challenging indoor datasets and showcase techniques for efficient mesh extraction. Specifically, we regularize the optimization procedure with depth information, enforce local smoothness of nearby Gaussians, and use off-the-shelf monocular networks to achieve better alignment with the true scene geometry. We propose an adaptive depth loss based on the gradient of color images, improving depth estimation and novel view synthesis results over various baselines. Our simple yet effective regularization technique enables direct mesh extraction from the Gaussian representation, yielding more physically accurate reconstructions of indoor scenes. Our code will be released in https://github.com/maturk/dn-splatter.
@@ 1 Introduction
Figure 1: Overview: We use depth and normal priors obtained from common handheld devices and general-purpose networks to enhance Gaussian splatting reconstruction quality. By regularizing Gaussian positions, local smoothness, and orientations, we demonstrate improvements in novel view synthesis and achieve more accurate mesh reconstructions on a variety of challenging indoor room datasets.
@@
The demand for high-fidelity 3D reconstruction of typical environments is increasing due to VR and AR applications. However, photorealistic and accurate 3D reconstruction of common indoor scenes from casually captured sensor data remains a persistent problem in 3D computer vision. Textureless and less-observed regions cause ambiguities in reconstructions and do not provide enough constraints for valid geometric solutions. Recently, neural implicit representations have achieved success in high-fidelity 3D reconstruction by representing scenes as a continuous volume with fully differentiable properties [31, 50, 25, 2]. However, the reconstruction of everyday indoor scenes still poses challenges, even for state-of-the-art methods. These methods rarely achieve good results in both photorealism and geometry reconstruction and often suffer from long training and rendering times, making them inaccessible for general use and VR/AR applications.
@@
3D Gaussian splatting [19] introduces a novel method for inverse rendering by representing a scene by many differentiable 3D Gaussian primitives with optimizable properties. This explicit representation enables real-time rendering of large, complex scenes–a capability that most neural implicit models lack. It results in a more interoperable scene representation since the scene appearance and geometry are directly expressed by the location, shape, and color attributes of Gaussians. However, due to the lack of 3D cues and surface constraints during optimization, artifacts and ambiguities are likely to occur, resulting in floaters and poor surface reconstruction. Scenes often contain millions of Gaussians, and their properties are directly modified by gradient descent based on photometric losses only. Little focus has been given to exploring better regularization techniques that result in visually and geometrically smoother and more plausible 3D reconstructions that can be converted into meshes, an important downstream application.
@@
Although many modern smartphones are equipped with low-resolution depth sensors, these are rarely used for novel view synthesis tasks. Motivated by this and advances in depth and normal estimation networks [3, 53, 7, 1], we explore the regularization of 3D Gaussian splatting with these geometric priors. Our goal is to enhance both photorealism and surface reconstruction in challenging indoor scenes. By designing an optimization strategy for 3D Gaussian splatting with depth and normal priors, we improve novel view synthesis results over baselines while respecting the captured scene geometry. We regularize the position of Gaussians with an edge-aware depth constraint and estimate normals from Gaussians to align them with the real surface boundaries estimated via monocular networks. We show how this simple regularization strategy, illustrated in Fig. 1, enables the extraction of meshes from the Gaussian scene representation, resulting in smoother and more geometrically accurate reconstructions. In summary, we make the following contributions:
@@
  • We use monocular normal priors to align Gaussians with the scene geometry and demonstrate how this aids reconstruction compared to alternative approaches.
  • We showcase how this regularization strategy enables efficient mesh extraction directly from the Gaussian scene.
@@ 2 Related work
Here we give an introduction to image-based rendering (IBR) methods for scene reconstruction and an overview of prior regularization strategies.
2.1 Traditional IBR
Reconstructing 3D geometry from images is a long-standing problem in computer vision. Traditional techniques enabled by Structure-from-Motion (SfM) [43, 40] and multi-view stereo (MVS) [9] techniques focused on reconstructing geometry as a sparse set of 3D points from images [21, 39]. The corresponding learning-based approaches [13, 30, 28, 32] usually replace some parts of the pipeline with differentiable modules, leading to a more accurate scene representation. Some work focuses on dense point reconstruction, normal estimation [41] or constructing triangle meshes from point sets [18, 27] to render novel views. However, traditional methods yield poor reconstruction results in large-scale, textureless scenes. In this paper, we demonstrate that incorporating geometric cues can significantly enhance scene reconstructions.
2.2 Neural implicit IBR
@@
The most success has been achieved with neural-based inverse rendering methods, most notably NeRF [31], which applies volume rendering [17] to represent scenes as continuous volumes with attributes encoded within a neural network. However, the 3D geometry extracted from these scenes is often ill-defined and suffers from artifacts and floaters. Subsequent work has focused on improving the rendering quality and scene reconstruction through regularization techniques and by adapting other scene representations such as signed distance functions [58] or occupancy grids.
@@
Prior regularization. Prior regularization of neural implicit models has been an active area of research. Previous NeRF-based approaches add depth regularization to explicitly supervise ray termination [6, 51] or impose smoothness constraints [34] on rendered depth maps. Other works explore regularizing with multi-view consistency [10, 6, 24] in sparse view settings. For SDF-based models, Manhattan-SDF [11] uses planar constraints on walls and flat surfaces to improve indoor reconstruction, and MonoSDF [58] uses depth and normal monocular estimates for scene geometry regularization. In this work, we investigate the regularization of 3D Gaussian splatting optimization with depth and normal priors to enhance photometric and geometric reconstruction.
@@
Meshable implicit representations. Surface extraction as triangle meshes is an important problem since most computer graphics pipelines still rely on triangle rasterization. Watertight meshes also provide a good approximation of scene geometry and surface quality, leading to the development of various metrics for mesh quality. Prior work has focused on extracting meshes from NeRF representations [37, 46, 45] with some success, but these methods often rely on expensive post-refinement stages. Most state-of-the-art techniques use SDF or occupancy representations [58, 54, 50, 25] combined with marching cubes [27] to achieve finer details. These methods involve querying and evaluating dense 3D volumes, often at multiple levels of detail, and are generally slow to train. In this work, we investigate extracting meshable surfaces directly from the explicit Gaussian scene representation.
@@
Meshable 3D Gaussians. Extracting meshable surfaces from Gaussian primitives is a relatively new topic. Keselman et al. [20] propose generating an oriented point set from a trained Gaussian scene to be meshed with Poisson reconstruction [18], using back-projected depth maps and analytically estimated normals. However, without regularization, this approach results in noisy point clouds that are difficult to mesh. NeuSG [4] addresses this by jointly training a dense SDF neural implicit model with the Gaussian scene, aligning Gaussians with SDF-estimated normals. While this method produces good reconstructions, it has long training times — over 16 hours on high-end GPUs — diminishing the appeal of 3DGS.
@@
SuGaR [12] proposes treating the positions of Gaussians and their densities as intersections of a level set and optimizes a signed-distance loss to converge Gaussians to the surface. Normals are estimated from the derivative of the signed distance, similar to Keselman et al. However, due to the lack of geometric priors, the reconstructions remain noisy. SuGaR refines the coarse mesh through differentiable optimization, making the process computationally costly. In contrast, the recent concurrent method 2DGS [16] proposes to use a Gaussian surfel representation to explicitly model planar surfaces. However, as shown in Table 1 and Table 2, without prior regularization, the results on outward-facing indoor datasets are poor. We demonstrate how to improve mesh extraction and novel view synthesis using depth and normal priors during optimization.
@@ 3 Preliminaries
Our work builds on 3D Gaussian splatting (3DGS, [19]) and we briefly describe the rasterization algorithm. 3DGS represents a scene with differentiable 3D Gaussian primitives parameterized by their mean $\bm{\mu}\in\mathbb{R}^{3}$ and covariance matrix $\bm{\Sigma}\in\mathbb{R}^{3\times 3}$ which is decomposed into a scaling vector $\bm{s}\in\mathbb{R}^{3}$ and a rotation quaternion $\bm{q}\in\mathbb{R}^{4}$. Other parameters include opacity $o\in\mathbb{R}$ and color $\bm{c}\in\mathbb{R}^{3}$, represented via spherical harmonics. Rendering a new view involves projecting 3D Gaussians into camera space as 2D Gaussians. These 2D Gaussians are sorted by z-depth in a single global sort and alpha-composited using the discrete volume rendering equation to produce pixel colors $\bf\hat{C}$:
@@
Gaussian projection and blending operations are parallelized, allowing for real-time rendering performance. The scene is initialized with Gaussian means and colors obtained from SfM [40, 41]. During optimization, the parameters of the Gaussians are updated via gradient descent through many rendering iterations to best fit the training dataset images. The algorithm progressively culls, splits, and duplicates Gaussians in the scene at fixed intervals based on Gaussian opacity, screen-space size, and the magnitude of the gradient of Gaussian means, respectively.
@@ 4 Method
In Section 4.1 we utilize sensor and monocular depth priors to regularize Gaussian positions with an edge-aware loss. Next, we extract normal directions from Gaussians and utilize these normal cues for regularization in Section 4.2. Additionally, we add a smoothing prior on rendered normal maps to better align nearby Gaussians during optimization in Section 4.3. Lastly, in Section 4.4, we use the optimized Gaussian scene to directly extract meshes using Poisson surface reconstruction.
@@ 4.1 Leveraging depth cues
where $\textit{d}_{i}$ is the $i^{\text{th}}$ Gaussian z-depth coordinate in view space. Since 3DGS does not sort Gaussians individually per-pixel along a viewing ray, and instead relies on a single global sort for efficiency; this is only an approximation of per-pixel depth. The z-depth ordering for nearby Gaussians and specific orientations is not guaranteed to be correct after 2D projection and a global sort, as explained in StopThePop[36]. However, sorting by depth for each pixel would be too computationally costly since scenes can contain millions of Gaussians. Nevertheless, this approximation remains effective, particularly for more regular geometries typically encountered in indoor datasets. We normalize depth estimates with the final accumulated transmittance $T_{i}$ per pixel, ensuring that we correctly estimate a depth value for pixels where the accumulated transmittance does not equal 1. We rasterize color and depths simultaneously per pixel in a single CUDA kernel forward pass, improving inference and training speed compared to separate rendering steps.
Sensor depth regularization. We directly apply depth regularization on predicted depth maps for datasets containing lidar or sensor depth measurements [38, 56, 44]. Common commercial depth sensors, especially low-resolution variants found in consumer devices like iPhones, often produce non-smooth edges at object boundaries and provide inaccurate readings. Based on this observation and inspired by [5, 22], we propose a gradient-aware depth loss for adaptive depth regularization based on the current RGB image. The depth loss is lowered in regions with large image gradients, signifying edges, ensuring that regularization is more enforced on smoother texture-less regions that typically pose challenges for photometric loss alone. Additionally, our experiments (cf. Table 5) show that using a logarithmic penalty results in smoother reconstructions compared to linear or quadratic penalties. This insight drives our formulation of the gradient-aware depth loss, which effectively balances the regularization across different regions of the image, adapting to the scene’s geometry and texture complexity. Specifically, the gradient-aware depth loss is defined as follows:
@@
$\hat{a},\hat{b}=\operatorname*{arg\,min}_{a,b}\sum_{ij}\|(a*D_{\text{mono},ij}+b)-D_{\text{sparse},ij}\|_{2}^{2}$ (5)
@@ 4.2 Leveraging normal cues
${\bm{\hat{n}}_{i}}=R\cdot\textrm{OneHot}{(\operatorname*{arg\,min}(s_{1},s_{2},s_{3}))}$ (6)
where OneHot(.) $\in\mathbb{R}^{3}$ returns a unit vector with zeros everywhere except at the position where the scaling $\bm{s}_{i}=(s_{1},s_{2},s_{3})$ is minimum. We minimize one of the scaling axes during training to force Gaussians to become disc-like surfels:
@@
In contrast to prior research [26, 8], which appends an additional learnable parameter per Gaussian for normal prediction, our approach derives normals directly from the geometry. Consequently, during back-propagation, adjustments to Gaussian scale and rotation parameters, i.e., covariance matrices, directly lead to updates in the normal estimates. Therefore, no additional learnable parameters are needed. Intuitively, this results in Gaussians better conforming to the scene’s geometry, as their orientations and scales are compelled to align with the surface normal.
@@
$\mathcal{L}_{\text{smooth}}=\frac{1}{|\hat{N}|}\sum_{i,j}\left(|\bm{\hat{N}}_{i+1,j}-\bm{\hat{N}}_{i,j}|+|\bm{\hat{N}}_{i,j+1}-\bm{\hat{N}}_{i,j}|\right)$ (10)
where $\hat{N}_{ij}$ represents estimated normal values at pixel position (i, j). Thus, our normal regularization loss is defined as follows: $\mathcal{L}_{\text{normal}}=\mathcal{L}_{\it{\hat{N}}}+\mathcal{L}_{\text{smooth}}$.
Normal initialization. We initialize Gaussian orientations by estimating normal directions from the initial SfM point cloud using [60], aligning the Gaussian orientations $q$ based on these estimated normals, and setting one of the scaling axes to be smaller than the others. This initialization helps to speed up convergence.

### v3

@@ Abstract
High-fidelity 3D reconstruction of common indoor scenes is crucial for VR and AR applications. 3D Gaussian splatting, a novel differentiable rendering technique, has achieved state-of-the-art novel view synthesis results with high rendering speeds and relatively low training times. However, its performance on scenes commonly seen in indoor datasets is poor due to the lack of geometric constraints during optimization. In this work, we explore the use of readily accessible geometric cues to enhance Gaussian splatting optimization in challenging, ill-posed, and textureless scenes. We extend 3D Gaussian splatting with depth and normal cues to tackle challenging indoor datasets and showcase techniques for efficient mesh extraction. Specifically, we regularize the optimization procedure with depth information, enforce local smoothness of nearby Gaussians, and use off-the-shelf monocular networks to achieve better alignment with the true scene geometry. We propose an adaptive depth loss based on the gradient of color images, improving depth estimation and novel view synthesis results over various baselines. Our simple yet effective regularization technique enables direct mesh extraction from the Gaussian representation, yielding more physically accurate reconstructions of indoor scenes. Our code will be released in https://github.com/maturk/dn-splatter.
11footnotetext: Denotes equal contribution
@@ 1 Introduction
Figure 1: Overview: We use depth and normal priors obtained from common handheld devices and general-purpose networks to enhance Gaussian splatting reconstruction quality. By regularizing Gaussian positions, local smoothness, and orientations, we demonstrate improvements in novel-view synthesis and achieve more accurate mesh reconstructions on a variety of challenging indoor room datasets.
@@
The demand for high-fidelity 3D reconstruction of typical environments is increasing due to VR and AR applications. However, photorealistic and accurate 3D reconstruction of common indoor scenes from casually captured sensor data remains a persistent challenge in 3D computer vision. Textureless and less-observed regions cause ambiguities in reconstructions and do not provide enough constraints for valid geometric solutions. Recently, neural implicit representations have achieved success in high-fidelity 3D reconstruction by representing scenes as continuous volumes with fully differentiable properties [30, 50, 25, 2]. However, the reconstruction of everyday indoor scenes still poses challenges, even for state-of-the-art methods. These methods rarely achieve good results in both photorealism and geometry reconstruction and often suffer from long training and rendering times, making them inaccessible for general use and VR/AR applications.
@@
3D Gaussian splatting [19] introduces a novel method for rendering by representing a scene by many differentiable 3D Gaussian primitives with optimizable properties. This explicit representation enables real-time rendering of large, complex scenes–a capability that most neural implicit models lack. 3DGS is a more interoperable scene representation compared to neural methods since it encodes a scene’s appearance and geometry directly through the location, shape, and color attributes of Gaussians. However, due to the lack of 3D cues and surface constraints during optimization, artifacts and ambiguities are likely to occur, resulting in floaters and poor surface reconstruction. Scenes can often contain millions of Gaussians, and their properties are directly modified by gradient descent based on photometric losses only. Little focus has been given to exploring better regularization techniques that result in visually and geometrically smoother and more plausible 3D reconstructions that can be converted into meshes, an important downstream application.
@@
Although many modern smartphones are equipped with low-resolution depth sensors, these are rarely used for novel-view synthesis tasks. Motivated by this and advances in depth and normal estimation networks [3, 53, 7, 1], we explore the regularization of 3D Gaussian splatting with these geometric priors. Our goal is to enhance both photorealism and surface reconstruction in challenging indoor scenes. By designing an optimization strategy for 3D Gaussian splatting with depth and normal priors, we improve novel-view synthesis results over baselines whilst better respecting the captured scene geometry. We regularize the position of Gaussians with an edge-aware depth constraint and estimate normals from Gaussians to align them with the real surface boundaries estimated via monocular networks. We show how this simple regularization strategy, illustrated in Fig. 1, enables the extraction of meshes from the Gaussian scene representation, resulting in smoother and more geometrically accurate reconstructions. In summary, we make the following contributions:
@@
  • We use monocular normal priors to align Gaussians with the scene and demonstrate how this aids reconstruction.
  • We show how regularization with depth and normal cues enables efficient mesh extraction directly from the Gaussian scene with improved novel-view synthesis.
@@ 2 Related work
Here we give a brief overview to image-based rendering (IBR) methods for scene reconstruction and an overview of prior methods utilizing geometry cues for regularization.
Traditional IBR. Reconstructing 3D geometry from images is a longstanding challenge in computer vision. Traditional methods like Structure-from-Motion (SfM) [43, 40] and Multi-view Stereo (MVS) [9, 60] techniques have focused on reconstructing geometry via a sparse set of multi-view consistent 3D points obtained by triangulation of features from images [21, 39]. Learning-based approaches [13, 29, 27, 31] usually replace parts of the pipeline with learnable modules, leading to improvements in the generalizability of the methods. Other work focus on normal estimation [41] and constructing triangle meshes [18, 26].
@@
Neural implicit IBR. Most success obtaining both photorealisic and geometrically accurate 3D reconstruction has been achieved with neural-based inverse rendering methods, most notably that of NeRF [30], which represents scenes as volumes with attributes encoded within a neural network and applies volume rendering [17] to achieve impressive novel-view results. However, the 3D geometry extracted from NeRFs are often poorly defined and suffer from artifacts and floaters. Subsequent work has focused on improving the rendering quality and scene reconstruction through regularization techniques [38, 6] or by adopting other scene representations such as signed distance functions [59, 50, 25] (SDFs) to improve geometry extraction.
@@
Prior regularization. Prior regularization of neural implicit models has been an active area of research. Previous NeRF-based approaches add depth regularization to explicitly supervise ray termination [6, 51, 38] or impose smoothness constraints [33] on rendered depth maps. Other works explore regularizing with multi-view consistency [10, 6, 24] in sparse view settings. For SDF-based models, Manhattan-SDF [12] uses planar constraints on walls and flat surfaces to improve indoor reconstruction, and MonoSDF [59] uses depth and normal monocular estimates for scene geometry regularization. In this work, we investigate the regularization of 3D Gaussian splatting optimization with depth and normal priors to enhance photometric and geometric reconstruction.
@@
Meshable implicit representations. Surface extraction as triangle meshes is an important problem since most computer graphics pipelines still rely on triangle rasterization. Watertight meshes also provide a good approximation of scene geometry and surface quality, leading to the development of various metrics for mesh quality. Prior work has focused on extracting meshes from NeRF representations [36, 46, 45] with some success, but these methods often rely on expensive post-refinement stages. Most State-of-the-Art techniques use SDF or occupancy representations [59, 54, 50, 25] combined with marching cubes [26] to achieve finer details. These methods involve querying and evaluating dense 3D volumes, often at multiple levels of detail, and are generally slow to train. In this work, we investigate extracting meshable surfaces directly from a Gaussian scene representation.
@@
Meshable 3D Gaussians. Extracting meshable surfaces from Gaussian primitives is a relatively new topic. Keselman et al. [20] propose generating an oriented point set from a trained Gaussian scene to be meshed with Poisson reconstruction [18], using back-projected depth maps and analytically estimated normals. However, without regularization, this approach results in noisy point clouds that are difficult to mesh. NeuSG [4] addresses this by jointly training a dense SDF neural implicit model [50] alongside a Gaussian scene, aligning Gaussians with SDF-estimated normals. The authors improve surface extraction from the Gaussian scene using SDF guidance; however, the approach requires long training times — over 16 hours on high-end GPUs — diminishing the appeal of 3DGS.
@@
SuGaR [11] proposes treating the positions of Gaussians as intersections of a level set and optimizes a signed-distance loss to converge Gaussians to the surface. Normals are estimated from the derivative of the signed distance, similar to Keselman et al. However, due to the lack of geometric priors, the reconstructions remain noisy (we show experimental comparing SuGaR meshing strategy in the supplement). SuGaR refines the coarse mesh with additional optimization, making the process computationally costly. In contrast, 2DGS [16] proposes to use a Gaussian surfel representation to explicitly model planar surfaces. However, we show in our experiments that without prior regularization, results on outward-facing indoor reconstructions are poor.
@@ 3 Preliminaries
Our work builds on 3D Gaussian splatting (3DGS, [19]) and we briefly describe the rasterization algorithm. 3DGS represents a scene using differentiable 3D Gaussian primitives, parameterized by their mean $\bm{\mu}\in\mathbb{R}^{3}$, a covariance matrix $\bm{\Sigma}\in\mathbb{R}^{3\times 3}$ decomposed into a scaling vector $\bm{s}\in\mathbb{R}^{3}$ and a rotation quaternion $\bm{q}\in\mathbb{R}^{4}$, along with opacity $o\in\mathbb{R}$ and color $\bm{c}\in\mathbb{R}^{3}$, represented via spherical harmonics. Rendering a new view involves projecting 3D Gaussians into 2D Gaussians in camera space. These 2D Gaussians are z depth-sorted and alpha-composited using the discrete volume rendering equation to produce a pixel color $\bf\hat{C}$:
@@
The scene is typically initialized with sparse SfM points obtained from a pre-processing step [40, 41]. In this work, we also explore initialization using sensor depth readings. The Gaussian scene is optimized using the Adaptive Density Control (ADC) algorithm [19], which progressively culls, splits, and duplicates Gaussians in the scene at fixed intervals based on Gaussian opacity, screen-space size, and the magnitude of the gradient of Gaussian means, respectively.
@@ 4 Method
We address the problem of achieving high-fidelity reconstruction of common indoor scenes that is both photorealistic and geometrically precise. In Section 4.1 we utilize sensor and monocular depth priors to regularize Gaussian positions with an edge-aware loss. Next, we extract normal directions from Gaussians and utilize normal cues for regularization in Section 4.2. Additionally, we add a smoothing prior on rendered normal maps to better align nearby Gaussians during optimization in Section 4.4 and initialize the Gaussian scene using dense depth information in Section 4.3. Lastly, in Section 4.5, we use the optimized Gaussian scene to directly extract meshes using Poisson surface reconstruction.
@@ 4.1 Leveraging depth cues
where $\textit{d}_{i}$ is the $i^{\text{th}}$ Gaussian z-depth coordinate in view space. Since 3DGS does not sort Gaussians individually per-pixel along a viewing ray, and instead relies on a single global sort for efficiency; this is only an approximation of per-pixel depth as explained in [35]. Nevertheless, this approximation remains effective, particularly for more regular geometries typically encountered in indoor datasets. We normalize depth estimates with the final accumulated transmittance $T_{i}$ per pixel, ensuring that we correctly estimate a depth value for pixels where the accumulated transmittance does not equal 1. We rasterize color and depths simultaneously per pixel in a single CUDA kernel forward pass, improving inference and training speed compared to separate rendering steps.
Sensor depth regularization. We directly apply depth regularization on predicted depth maps for datasets containing lidar or sensor depth measurements [37, 57, 44]. Common commercial depth sensors, especially low-resolution variants found in consumer devices like iPhones, often produce non-smooth edges at object boundaries and provide inaccurate readings. Based on this observation and inspired by [5, 22], we propose a gradient-aware depth loss for adaptive depth regularization based on the current RGB image. The depth loss is lowered in regions with large image gradients, signifying edges, ensuring that regularization is more enforced on smoother texture-less regions that typically pose challenges for photometric regularization alone. Additionally, our experiments (cf. Table 6) show that using a logarithmic penalty results in smoother reconstructions compared to linear or quadratic penalties. This insight drives our formulation of the gradient-aware depth loss, which effectively balances the regularization across different regions of the image, adapting to the scene’s geometry and texture complexity. We define the gradient-aware depth loss as follows:
@@
$\hat{a},\hat{b}=\operatorname*{arg\,min}_{a,b}\sum_{ij}\|(a*D_{\text{mono},ij}+b)-D_{\text{sparse},ij}\|_{2}^{2},$ (5)
@@ 4.2 Leveraging normal cues
${\bm{\hat{n}}_{i}}=R\cdot\textrm{OneHot}{(\operatorname*{arg\,min}(s_{1},s_{2},s_{3}))},$ (6)
where $\text{OneHot}(.)\in\mathbb{R}^{3}$ returns a unit vector with zeros everywhere except at the position where the scaling $\bm{s}_{i}=(s_{1},s_{2},s_{3})$ is minimum. $R$ is obtained from the quaternion $q=(w,x,y,z)^{\top}$ using:
$R=\begin{bmatrix}1-2\left(y^{2}+z^{2}\right)&2\left(xy-wz\right)&2\left(xz+wy\right)\\ 2\left(xy+wz\right)&1-2\left(x^{2}+z^{2}\right)&2\left(yz-wx\right)\\ 2\left(xz-wy\right)&2\left(yz+wx\right)&1-2\left(x^{2}+y^{2}\right)\\ \end{bmatrix}$ (7)
We minimize one of the scaling axes during training to force Gaussians to become disc-like surfels:
@@
This approach derives normal estimates directly from the geometry of Gaussians. Consequently, during back-propagation, adjustments to scale and rotation parameters, i.e., covariance matrices, directly lead to updates in normal estimates. Therefore, no additional learnable parameters are needed. Intuitively, this results in Gaussians better conforming to the scene’s geometry, as their orientations and scales are compelled to align with the surface normal.
@@
$\mathcal{L}_{\text{smooth}}=\sum_{i,j}\left(|\nabla_{i}\bm{\hat{N}}_{i,j}|+|\nabla_{j}\bm{\hat{N}}_{i,j}|\right),$ (11)
where $\hat{N}_{ij}$ represents estimated normal values at pixel position (i, j) and $\nabla$ represents the finite difference operator that convolves its input with $[-1,1]$ for the $i$-axis and $[-1,1]^{\top}$ for the $j$-axis. Thus, our normal regularization loss is defined as $\mathcal{L}_{\text{normal}}=\mathcal{L}_{\it{\hat{N}}}+\mathcal{L}_{\text{smooth}}.$
@@
4.3 Gaussian initialization
Rather than relying on SfM points for initialization, we make use of sensor depth readings, where available, by back-projecting depths into world coordinates for dense scene initialization. Additionally, we initialize Gaussian orientations by estimating normal directions from the initial point cloud using [61], aligning the Gaussian orientations $q$ with these estimated normals using Eq. 7, and setting one of the scaling axes smaller than the others. This initialization helps normal estimation convergence.